# Minimax estimation of functionals in sparse vector model with correlated observations

### v1

@@ 1 Introduction
$Q({\boldsymbol{\theta}})=\sum_{i=1}^{d}\theta_{i}^{2},\quad\|{\boldsymbol{\theta}}\|_{2}=\sqrt{Q({\boldsymbol{\theta}})}.$
@@
Most of the existing literature on estimation of functionals and on the related problem of testing under sparse vector models is focused on the case of uncorrelated data (${\boldsymbol{\Sigma}}$ is a diagonal matrix). However, in many applications such as time series analysis one needs to deal with dependent observations. In this paper, we study the setting where only partial information on ${\boldsymbol{\Sigma}}$ is available, namely, either the Frobenius norm of ${\boldsymbol{\Sigma}}$ or an upper bound is known. We propose estimators of $Q({\boldsymbol{\theta}})$ and $\|{\boldsymbol{\theta}}\|_{2}$ that are adaptive to the noise level and rate optimal in a minimax sense for $\|{\boldsymbol{\theta}}\|_{2}$. As a consequence, we construct minimax optimal tests for signal detection. Furthermore, we propose minimax optimal estimators of the noise level $\sigma$.
@@
Carpentier et al. [4, 5] further extended these results to the model of high-dimensional sparse regression with i.i.d. noise. Model (1) in the case of diagonal matrix $\bf{\Sigma}$ with known entries and $\sigma=1$ was analyzed in Laurent et al. [14], Chhor et al. [6]. Under such particular conditions, Laurent et al. [14] studied the minimax rates for the problem of testing mentioned above. In Chhor et al. [6] these results are extended to testing of $H_{0}:{\boldsymbol{\theta}}=\bf{0}$ against sparse alternatives separated from $\bf{0}$ in the $\ell_{p}$-norm with $1\leq p<\infty$ and further to estimation of the corresponding norms. Another recent development due to Cannone and George [3] shows that introducing an arbitrary contamination in model (1) with ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ leads to slower minimax rates (for the problem of testing with $\ell_{2}$-separation) than those given in [7]. Finally, to the best of our knowledge, the only related works allowing for correlated noise, again in the area of minimax rates of testing, are Liu et al. [15] and Kotekal and Gao [12]. The paper [15] deals a change-point framework and establishes the minimax rates of testing for the particular case $s=d$ (no sparsity) assuming that matrix ${\boldsymbol{\Sigma}}$ is fixed and positive definite. In [12] the authors consider model (1) with $\sigma=1$, equicorrelation matrix ${\boldsymbol{\Sigma}}=(1-\gamma){\boldsymbol{I}}+\gamma{\boldsymbol{1}}{\boldsymbol{1}}^{\top}$, where $\gamma>0$, and ${\boldsymbol{1}}$ is a $d$-dimensional column vector with all entries equal to $1$, and establish the minimax rates of testing the hypothesis $H_{0}:{\boldsymbol{\theta}}=\bf{0}$ against sparse alternatives separated from $\bf{0}$ in the $\ell_{2}$-norm assuming that $\gamma$ is known. In a more recent work [13], the same authors find the minimax rates of estimating linear functionals under the equicorrelated Gaussian noise.
@@
In summary, the existing literature (mostly focused on the problem of testing rather than estimation of functionals) allows one to tackle the problem under sparsity when ${\boldsymbol{\Sigma}}$ is either a diagonal matrix with known entries or the equicorrelation matrix, and in the latter case all parameters of the model except ${\boldsymbol{\theta}}$ need to be known. However, under general correlation structure, the question of how to estimate functionals of a sparse vector ${\boldsymbol{\theta}}$ in a minimax optimal way and how to construct the associated tests remains open. In this paper, we focus on the problem of estimating the functionals $Q({\boldsymbol{\theta}})$ and $\|{\boldsymbol{\theta}}\|_{2}$ and we consider the general setting, with arbitrary structure of ${\boldsymbol{\Sigma}}$ and with the noise level $\sigma$ that can be either known or unknown. Moreover, the estimators that we propose do not require knowing the off-diagonal entries of ${\boldsymbol{\Sigma}}$. They need only partial information about ${\boldsymbol{\Sigma}}$ such as the relative scale of the diagonal entries, the value of its Frobenius norm or an upper bound on it. We construct estimators that are optimal in a minimax sense and are adaptive to the noise level. We also derive the associated minimax rates of testing. The minimax rates that we obtain for estimating the $\ell_{2}$-norm under known and unknown $\sigma$ are of the form (2) and (3), respectively, with $t=\|{\boldsymbol{\Sigma}}\|_{F}^{2}$, where $\|{\boldsymbol{\Sigma}}\|_{F}$ is the Frobenius norm of ${\boldsymbol{\Sigma}}$. If ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ we recover the rates $\sigma^{2}\phi(s,d)$ and $\sigma^{2}\phi^{*}(s,d)$ obtained in [7, 8]. For general ${\boldsymbol{\Sigma}}$, we arrive at an interesting conclusion that the possibility of faster or slower rates is leveraged not by the dimension $d$ but by the value of the Frobenius norm of ${\boldsymbol{\Sigma}}$.
@@
Obtaining the results for the model with correlated noise is challenging. The tools used in [8] for the setting with unknown $\sigma$, including exponential inequalities and sample splitting to obtain multiple independent estimators, break down when the observations are correlated. We develop new proof techniques allowing us to obtain an improvement upon [8] even in the case of i.i.d. observations since our estimator does not need sample splitting.
@@
Notation Let $[n]:=\{1,\dots,n\}$. For $1\leq q\leq\infty$, let $\|\cdot\|_{q}$ denote the $\ell_{q}$-norm in $\mathbb{R}^{d}$, and let $B_{q}(r):=\{x\in\mathbb{R}^{d}:\|x\|_{q}\leq r\}$ be the $\ell_{q}$-ball of radius $r$. We denote by $\Phi$ the cumulative distribution function of the standard Gaussian distribution. For $x,y\in\mathbb{R}$, we set $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. We use standard big-$O$ notations, e.g., for two positive sequence $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some absolute constant $C>0$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$ if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$. Given a matrix $A\in\mathbb{R}^{d\times d}$ and $\mathcal{I},\mathcal{J}\subseteq[d]$, we denote by $A_{\mathcal{I},\mathcal{J}}$ the submatrix of $A$ with rows from $\mathcal{I}$ columns from $\mathcal{J}$, by $A_{ij}$ the entries of $A$, and by $\|A\|_{F}$, $\lambda_{\max}(A)$, and $\mathbf{tr}[A]$, the Frobenius norm, the maximal eigenvalue, and the trace of $A$, respectively. We denote by ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}$ the expectation with respect to the probability measure ${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}$ of random vector ${\boldsymbol{Y}}$ satisfying (1). We denote by $c,C,C^{\prime}$ absolute positive constants that can vary from line to line.
@@ 2 Estimating the quadratic functional and the $\ell_{2}$-norm when $\sigma$ is known
Assumption LABEL:as:var0 is just a normalization condition. It assures that $\sigma$ can be considered as the noise level. As ${\boldsymbol{\Sigma}}$ is a covariance matrix Assumption LABEL:as:var0 implies that all entries of ${\boldsymbol{\Sigma}}$ are bounded by 1. Note also that we can replace the condition $\max_{1\leq i\leq d}\sigma_{i}^{2}\leq 1$ by $\max_{1\leq i\leq d}\sigma_{i}^{2}\leq c$, where $c>0$ is an absolute constant, with no incidence on the results.
@@
$\psi(s,{\boldsymbol{\Sigma}}):=\phi(s,\|{\boldsymbol{\Sigma}}\|_{F}^{2})=\left\{\begin{array}[]{lc}s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})&\textrm{if}\;s\leq\|{\boldsymbol{\Sigma}}\|_{F},\\ \|{\boldsymbol{\Sigma}}\|_{F}&\textrm{otherwise,}\end{array}\right.$
@@
$\bar{\psi}(s,{\boldsymbol{\Sigma}}):=\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s.$
@@
There exists an absolute constant $C>0$ such that, for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any ${\boldsymbol{\Sigma}}$ satisfying Assumption LABEL:as:var0,
@@
Furthermore, for any ${\boldsymbol{\Sigma}}$ satisfying Assumption LABEL:as:var0,
@@
It is interesting to compare (7) to the bound for the case ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ [7], that is, $\sigma^{2}\phi(s,d)$, where $\phi(s,d)$ is defined in (2). Since ${\boldsymbol{\Sigma}}$ is a covariance matrix, Assumption LABEL:as:var0 implies that $\|{\boldsymbol{\Sigma}}\|_{F}$ is at most of order $d$. Particularly, in the extreme case $\|{\boldsymbol{\Sigma}}\|_{F}\asymp d$, there is no elbow in the rate and $\psi(s,{\boldsymbol{\Sigma}})\asymp s\log(1+d^{2}/s^{2})$ for all $1\leq s\leq d$. Thus, we get a substantial deterioration in the zone $s>\sqrt{d}$ compared to the rate $\sigma^{2}\phi(s,d)$ with uncorrelated data. On the contrary, for the other extreme case $\|{\boldsymbol{\Sigma}}\|_{F}\asymp\sqrt{d}$, we have $\psi(s,{\boldsymbol{\Sigma}})\asymp\phi(s,d)$ and we recover the rate obtained in [7] for ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$. Finally, if ${\boldsymbol{\Sigma}}$ is a very sparse matrix, with very small diagonal elements (or some diagonal elements equal to 0, that is, no noise in some components of ${\boldsymbol{\theta}}$), then the rate can be further improved.
Moreover, the rate $\sigma^{2}\psi(s,{\boldsymbol{\Sigma}})$ and the estimator $\hat{N}$ are minimax optimal in the following sense. Consider the class of parameters
$\Theta(s,d):=\{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\mathbb{R}^{d}\times\mathbb{R}^{d\times d}:{\boldsymbol{\theta}}\in B_{0}(s,d)\;\textrm{and}\;{\boldsymbol{\Sigma}}_{ii}=1\ \textrm{for}\ i=1,\dots,d\}.$
We have the following minimax lower bound.
Theorem 2.
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big{(}\sigma^{-1}\psi(s,{\boldsymbol{\Sigma}})^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big{)}\geq c,$
where $\inf_{\hat{T}}$ denotes the infimum over all estimators.
Combining (7) and Theorem 2 with $w(u)=u^{2}$ we obtain the next corollary.
Corollary 1.
We have
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}\frac{{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[(\hat{T}-\|{\boldsymbol{\theta}}\|_{2})^{2}]}{\psi(s,{\boldsymbol{\Sigma}})}\asymp\sigma^{2}.$
Theorem 2 and Corollary 1 deal with scaled minimax mean square risk. The estimator $\hat{N}$, which is rate optimal with respect to such a risk, requires the knowledge of the Frobenius norm $\|{\boldsymbol{\Sigma}}\|_{F}$. If only an upper bound on $\|{\boldsymbol{\Sigma}}\|_{F}$ is available, we can modify the definition of $\hat{N}$ by using this upper bound in place of $\|{\boldsymbol{\Sigma}}\|_{F}$. Such an estimator is also minimax optimal, in a slightly different sense. Indeed, consider the class of pairs $({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})$ defined as
$\Theta(s,d,\rho):=\Theta(s,d)\cap\{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}}):\|{\boldsymbol{\Sigma}}\|_{F}\leq\rho\},$
@@
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big{(}\sigma^{-1}(\phi(s,\rho^{2}))^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big{)}\geq c,$

### v2

@@ 1 Introduction
$Q({\boldsymbol{\theta}})=\sum_{i=1}^{d}\theta_{i}^{2},\quad\|{\boldsymbol{\theta}}\|_{2}=\sqrt{Q({\boldsymbol{\theta}})}.$ (3)
@@
Most of the existing literature on estimation of functionals and on the related problem of testing under sparse vector models is focused on the case of uncorrelated data ($\boldsymbol{\Sigma}$} is a diagonal matrix). However, in many applications such as time series analysis one needs to deal with dependent observations. In this paper, we study the setting where only partial information on $\boldsymbol{\Sigma}$} is available, namely, either the Frobenius norm of $\boldsymbol{\Sigma}$} or an upper bound is known. We propose estimators of $Q({\boldsymbol{\theta}})$ and $\|{\boldsymbol{\theta}}\|_{2}$ that are adaptive to the noise level and are rate optimal in a minimax sense for $\|{\boldsymbol{\theta}}\|_{2}$. As a consequence, we construct minimax optimal tests for signal detection. Furthermore, we propose minimax optimal estimators of the noise level $\sigma$.
@@
Carpentier et al. [4, 5] further extended these results to the model of high-dimensional sparse regression with i.i.d. noise. Model (2) in the case of diagonal matrix $\bf{\Sigma}$ with known entries and $\sigma=1$ was analyzed in Laurent et al. [15], Chhor et al. [6]. Under such particular conditions, Laurent et al. [15] studied the minimax rates for the problem of testing mentioned above. In Chhor et al. [6] these results are extended to testing of $H_{0}:{\boldsymbol{\theta}}=\bf{0}$ against sparse alternatives separated from $\bf{0}$ in the $\ell_{p}$-norm with $1\leq p<\infty$ and further to estimation of the corresponding norms. Another recent development due to Cannone and George [3] shows that introducing an arbitrary contamination in model (2) with ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ leads to slower minimax rates (for the problem of testing with $\ell_{2}$-separation) than those given in [7]. Finally, there exist works allowing for correlated noise, again focused on the problem of testing, see Hall and Jin [10], Liu et al. [16], Kotekal and Gao [13]. The paper [16] deals with a change-point framework and establishes the minimax rates of testing for the particular case $s=d$ (no sparsity) assuming that matrix $\boldsymbol{\Sigma}$} is fixed and positive definite. In [13] the authors consider model (2) with $\sigma=1$, equicorrelation matrix ${\boldsymbol{\Sigma}}=(1-\gamma){\boldsymbol{I}}+\gamma{\boldsymbol{1}}{\boldsymbol{1}}^{\top}$, where $\gamma>0$, and $\boldsymbol{1}$} is a $d$-dimensional column vector with all entries equal to $1$, and establish the minimax rates of testing the hypothesis $H_{0}:{\boldsymbol{\theta}}=\bf{0}$ against sparse alternatives separated from $\bf{0}$ in the $\ell_{2}$-norm assuming that $\gamma$ is known. They show that, for some choices of $\gamma$, the detection boundary can be smaller than its value when the noise is uncorrelated. In a more recent work [14], the same authors find the minimax rates of estimating linear functionals under the equicorrelated Gaussian noise. A different asymptotic instance of the same testing problem is considered in Hall and Jin [10], where it is assumed that all non-zero components of $\theta$ have the same value $r\sqrt{\log d}$ for some $r>0$, the sparsity parameter satisfies $s=d^{a}$ for some $a\in(0,1/2)$, and the positions of non-zero components are randomly generated. In this model, it is proved in Hall and Jin [10] that if $\boldsymbol{\Sigma}$} is a Toeplitz matrix, then the asymptotic (as $d\to\infty$) detection boundary can be smaller than its value when the noise is uncorrelated. The improvement is not in the rate but in the constant factor.
@@
In summary, the existing literature dealing with our non-asymptotic setting (mostly focused on the problem of testing rather than estimation of functionals) allows one to tackle the problem under sparsity when $\boldsymbol{\Sigma}$} is either a diagonal matrix with known entries or the equicorrelation matrix, and in the latter case all parameters of the model except $\boldsymbol{\theta}$} need to be known. However, under general correlation structure, the question of how to estimate functionals of a sparse vector $\boldsymbol{\theta}$} in a minimax optimal way and how to construct the associated tests remains open. In this paper, we focus on the problem of estimating the functionals $Q({\boldsymbol{\theta}})$ and $\|{\boldsymbol{\theta}}\|_{2}$ and we consider the general setting, with arbitrary structure of $\boldsymbol{\Sigma}$} and with the noise level $\sigma$ that can be either known or unknown. Moreover, the estimators that we propose do not require knowing the off-diagonal entries of $\boldsymbol{\Sigma}$}. They need only partial information about $\boldsymbol{\Sigma}$} such as the relative scale of the diagonal entries, the value of its Frobenius norm or an upper bound on it. We construct estimators of $\|{\boldsymbol{\theta}}\|_{2}$ that are optimal in a minimax sense and are adaptive to the noise level $\sigma$. We also derive the associated minimax rates of testing. The minimax rates that we obtain for estimating $\|{\boldsymbol{\theta}}\|_{2}$ under known and unknown $\sigma$ are of the form (4) and (5), respectively, with $t=\|{\boldsymbol{\Sigma}}\|_{F}^{2}$, where $\|{\boldsymbol{\Sigma}}\|_{F}$ is the Frobenius norm of $\boldsymbol{\Sigma}$}. If ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ we recover the rates $\sigma^{2}\phi(s,d)$ and $\sigma^{2}\phi^{*}(s,d)$ obtained in [7, 8]. For general $\boldsymbol{\Sigma}$}, we arrive at an interesting conclusion that the possibility of faster or slower rates is leveraged not by the dimension $d$ but by the value of the Frobenius norm of $\boldsymbol{\Sigma}$}.
@@
Obtaining the results for the model with correlated noise is challenging. The tools used in [8] for the setting with unknown $\sigma$, including exponential inequalities and sample splitting to obtain multiple independent estimators, break down when the observations are correlated. We develop new proof techniques allowing us to obtain a slight improvement upon [8] even in the case of i.i.d. observations, which consists in the fact that our estimator does not need sample splitting.
This paper is organized as follows. In Section 2, we discuss estimation of $Q(\boldsymbol{\theta})$ and $\|\boldsymbol{\theta}\|_{2}$ when $\sigma$ is known. In Section 3, we introduce an estimator of $\sigma$. Section 4 deals with the problem of estimating $\|\boldsymbol{\theta}\|_{2}$ when $\sigma$ is unknown. In Section 5, we consider testing of the null hypothesis $\|\boldsymbol{\theta}\|_{2}=0$. Section 6 is devoted to the proofs.
@@
Notation Let $n]:=\{1,\dots,n\}$. For $1\leq q\leq\infty$, let $\|\cdot\|_{q}$ denote the $\ell_{q}$-norm in $\mathbb{R}^{d}$, and let $B_{q}(r):=\{x\in\mathbb{R}^{d}:\|x\|_{q}\leq r\}$ be the $\ell_{q}$-ball of radius $r$. We denote by $\Phi$ the cumulative distribution function of the standard Gaussian distribution. For $x,y\in\mathbb{R}$, we set $x\wedge y:=\min\{x,y\}$ and $x\vee y:=\max\{x,y\}$. We use standard big-$O$ notations, e.g., for two positive sequence $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$ if $a_{n}\leq Cb_{n}$ for some absolute constant $C>0$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$ if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$. Given a matrix $A\in\mathbb{R}^{d\times d}$ and $\mathcal{I},\mathcal{J}\subseteq[d]$, we denote by $A_{\mathcal{I},\mathcal{J}}$ the submatrix of $A$ with rows from $\mathcal{I}$ columns from $\mathcal{J}$, by $A_{ij}$ the entries of $A$, and by $\|A\|_{F}$, $\lambda_{\max}(A)$, and $\mathbf{tr}[A]$, the Frobenius norm, the maximal eigenvalue, and the trace of $A$, respectively. We denote by ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}$ the expectation with respect to the probability measure ${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}$ of random vector $\boldsymbol{Y}$} satisfying (2). We denote by $c,C,C^{\prime}$ absolute positive constants that can vary from line to line.
@@ 2 Estimating the quadratic functional and the $\ell_{2}$-norm when $\sigma$ is known
Assumption 1 is just a normalization condition. It assures that $\sigma$ can be considered as the noise level. As $\boldsymbol{\Sigma}$} is a covariance matrix Assumption 1 implies that all entries of $\boldsymbol{\Sigma}$} are bounded by 1. Note also that we can replace the condition $\max_{1\leq i\leq d}\sigma_{i}^{2}\leq 1$ by $\max_{1\leq i\leq d}\sigma_{i}^{2}\leq c$, where $c>0$ is an absolute constant, with no incidence on the results.
@@
$\psi(s,{\boldsymbol{\Sigma}}):=\phi(s,\|{\boldsymbol{\Sigma}}\|_{F}^{2})=\left\{\begin{array}[]{lc}s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})&\textrm{if}\;s\leq\|{\boldsymbol{\Sigma}}\|_{F},\\ \|{\boldsymbol{\Sigma}}\|_{F}&\textrm{otherwise,}\end{array}\right$. (8)
@@
$\bar{\psi}(s,{\boldsymbol{\Sigma}}):=\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s.$ (9)
@@
There exists an absolute constant $C>0$ such that, for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any $\boldsymbol{\Sigma}$} satisfying Assumption 1,
@@
Furthermore, for any $\boldsymbol{\Sigma}$} satisfying Assumption 1,
@@
It is interesting to compare (12) to the bound for the case ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$ [7], that is, $\sigma^{2}\phi(s,d)$, where $\phi(s,d)$ is defined in (4). Since $\boldsymbol{\Sigma}$} is a covariance matrix, Assumption 1 implies that $\|{\boldsymbol{\Sigma}}\|_{F}$ is at most of order $d$. Particularly, in the extreme case $\|{\boldsymbol{\Sigma}}\|_{F}\asymp d$, there is no elbow in the rate and $\psi(s,{\boldsymbol{\Sigma}})\asymp s\log(1+d^{2}/s^{2})$ for all $1\leq s\leq d$. Thus, we get a substantial deterioration in the zone $s>\sqrt{d}$ compared to the rate $\sigma^{2}\phi(s,d)$ with uncorrelated data. On the contrary, for the other extreme case $\|{\boldsymbol{\Sigma}}\|_{F}\asymp\sqrt{d}$, we have $\psi(s,{\boldsymbol{\Sigma}})\asymp\phi(s,d)$ and we recover the rate obtained in [7] for ${\boldsymbol{\Sigma}}={\boldsymbol{I}}$. Finally, if $\boldsymbol{\Sigma}$} is a very sparse matrix, with very small diagonal elements (or some diagonal elements equal to 0, that is, no noise in some components of $\boldsymbol{\theta}$}), then the rate can be further improved.
If only an upper bound on $\|{\boldsymbol{\Sigma}}\|_{F}$ is available, we can modify the definition of $\hat{N}$ by using this upper bound in place of $\|{\boldsymbol{\Sigma}}\|_{F}$. We now show that such an estimator is minimax optimal in the following sense. Consider the class of parameters
$\Theta(s,d,\rho):=\{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\mathbb{R}^{d}\times\mathbb{R}^{d\times d}:{\boldsymbol{\theta}}\in B_{0}(s,d),\;\|{\boldsymbol{\Sigma}}\|_{F}\leq\rho\;\textrm{and}\;{\boldsymbol{\Sigma}}_{ii}=1\ \textrm{for}\ i=1,\dots,d\},$ (13)
@@
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big(\sigma^{-1}(\phi(s,\rho^{2}))^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big)\geq c,$ (14)